# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v2

@@ Abstract
This paper proposes a novel approach for statistical modelling of a continuous random variable $X$ on $[0,1)$, based on its digit representation $X=.X_{1}X_{2}\ldots$. In general, $X$ can be coupled with a latent random variable $N$ so that $(X_{1},\ldots,X_{N})$ becomes a sufficient statistics and $X_{N+1}X_{N+2}\ldots$ is uniformly distributed. In line with this fact, and focusing on binary digits for simplicity, we propose a family of generalized finite Pólya trees that induces a random density for a sample, which becomes a flexible tool for density estimation. Here, the digit system may be random and learned from the data. We provide a detailed Bayesian analysis, including closed form expression for the posterior distribution which sidesteps the need of MCMC methods for posterior inference. We analyse the frequentist properties as the sample size increases, and provide sufficient conditions for consistency of the posterior distributions of the random density and $N$. We consider an extension to data spanning multiple orders of magnitude, and propose a prior distribution that encodes the so-called extended Newcomb-Benford law. Such a model shows promising results for density estimation of human-activity data. Our methodology is illustrated on several synthetic and real datasets.
@@ 1.3 Outline
The Supplemental material collects the missing proofs in the present paper.
@@ 6 Numerical illustrations
This section investigates several aspects of our models via numerical simulations. Prior elicitation follows the discussion in Section 4.3. Unless otherwise specified, a priori $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,20\}$ where the Poisson distribution has mean 5.
@@ 6.1 Posterior results when using GFPT1 and Pólya tree priors
Figures 46 show the mean across the 100 independent datasets of
@@
respectively. As expected Figures 46 show that larger values of $\alpha_{0}$ correspond to smaller values of $N$ (a posteriori) and vice versa. This is particularly evident for datasets DG1 and DG3. In general, we see that GFPT1 priors outperform PT priors in terms of the quality of the density estimate. Moreover, larger $\alpha_{0}$ clearly yield less wiggly functions when using PT priors, while this dependence is not so clear when using GFPT1 priors. In conclusion, we find that $\alpha_{0}=0.1$ provides the best trade off between wigglyness and accuracy of the density estimate, and we suggest using it as default value for GFPT1 priors.
Figure 4: Posterior results for the second simulation study in Section 6.1: Total variation (TV) distance between true and estimated densities as a function of the sample size $m$ where the three columns correspond to the data examples DG1–DG3, respectively. Figures are averaged over 100 independent replicates. Solid and dashed lines refer to the use of GFPT1 and PT priors, respectively. Different colors refer to different values of $\alpha_{0}$.
Figure 5: Posterior results for the second simulation study in Section 6.1: $W$-function for the difference between true and estimated densities. All details are as in the caption of Figure 4.
Figure 6: Posterior results for the second simulation study in Section 6.1: Posterior expectation of $N$ for the simulation in Section 6.1. All details are as in the caption of Figure 4.
@@ 6.3 Fitted models for real datasets spanning multiple orders of magnitude
Data for the first four dataset can be found at https://github.com/jasonlong/benfords-law, while data for the income dataset is publicly available from https://www.census.gov/data.html. Table 1 reports, for each dataset, the empirical range (maximum and minimum value of the data) and two diagnostics computed from the log-mantissa $U=\log_{10}(Z)-\lfloor\log_{10}(Z)\rfloor$: The Wasserstein-1 distance $W_{1}$ between the empirical distribution of $U$ and the $\mathrm{Unif}[0,1)$ law, and the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ to $\mathrm{Unif}[0,1)$. Under the Newcomb-Benford law, the distribution is scale invariant, i.e., the log-mantissa $U$ is $\mathrm{Unif}[0,1)$-distributed, hence, small values of $W_{1}$ and $d_{\mathrm{KS}}$ indicate that our MBPT model is well supported by the data. The datasets in Table 1 span four, three, eight, six, and seven orders of magnitude, respectively, making them particularly well suited to test the adequacy of the base-2 and base-10 multiscale Benford Pólya tree priors. In addition, the $W_{1}$ and $d_{\mathrm{KS}}$ statistics quantify how close each dataset is to the Benford benchmark: eurodist has the largest deviations (0.18 and 0.07), signalling weak scale-invariance; by contrast, census is extremely close to uniform (0.001 and 0.003), with income, GDP, and Twitter also showing relatively small deviations. This pattern anticipates our empirical findings, whereby the MBPT model delivers the best fit where the data spans a large number of orders of magnitude and $W_{1}$ and $d_{\mathrm{KS}}$ are close to zero.
@@
The posterior inferences using the different priors are compared in term of the widely applicable information criterion (WAIC), see Watanabe (2013), which is a consistent estimator of the out-of-sample error (Vehtari et al., 2017). Since the absolute value of the WAIC is irrelevant, we fix the DPM as a baseline prior and report in Table 2 the relative improvement of the WAIC for any model over the DPM (larger values of the relative improvement correspond to better model performance, while negative values indicate that the DPM achieves a better fit to the data). For datasets eurodist, GDP, and census, all priors seem to perform similarly. For the datasets Twitter and income datasets, the multiscale Benford Pólya tree priors clearly outperforms all other priors, where in particular, the $\mathrm{MBPT}_{10}$ prior shows a remarkable improvement of rougly 19% and 25% over the DPM prior, respectively. These results confirm the intuition obtained by looking at the diagnostics based on the log-mantissa and the span of the datasets discussed above. Moreover, in all scenarios, the WAIC when using the GFPT2 prior slightly improves as compared to the GFPT1 prior. Although not reported here, we have also tried fitting a variant of the MBPT models where the parameters $\alpha$ are not specified in accordance to the Newcomb-Benford law but according to the standard choice of (16). We noticed substantial deterioration in the WAICs across all datasets, and in particular for Twitter, for which this variant performs worse than the DPM baseline, and income, for which this variant performs on par to the baseline.
@@
Finally, we comment on the number of sufficient digits needed in the different datasets according to the GFPT and MBPT models. Table 3 reports the (estimated) a posteriori expectation of $N$ in the different models. Note that $N$ is typically small a posteriori, despite $n_{\max}$ being equal to six, four, seventeen, eight, and six in the different datasets respectively. Recall that the prior for $N$ is trucated at $n_{\max}$ for the $\mathrm{MBPT}_{10}$ prior and to $\lceil\log_{2}(10^{n_{\max}})\rceil$ for the other priors. Hence, the sufficient number of digits are effectively always less than $n_{\max}$. Except for the Twitter and income datasets, we observe that $N$ tends to be smaller a posteriori under the $\mathrm{MBPT}_{2}$ prior than under the GFPT1 and GFPT2 priors. This entails that the max-scaling of the data disaligns the data from the nested binary partitions. Indeed, using the GFPT2 prior, the posterior adaptively learns the NBP, leading to a smaller $N$ than under the GFPT1 prior that works with the fixed NBP. For all the different datasets, a posteriori $N$ tends to be smaller under the $\mathrm{MBPT}_{10}$ prior, however, the possible values taken by $X_{1:N}$ is $10^{N}$ under the $\mathrm{MBPT}_{10}$ prior and (approximately) $2^{N}$ under the other models. We report the corresponding (estimated) posterior expectations in Table 3. When using the $\mathrm{MBPT}_{10}$ prior a larger number of possible digits is required a posteriori than if using another of the priors, but compared with the $\mathrm{MBPT}_{2}$ prior, the difference may not be extreme, since the scale is coarser for $10^{N}$ than for $2^{N}$. On the other hand, there are several situations in which the $\mathrm{GPFT2}$ prior causes a posteriori a much lower number of possible digits than if using another of the priors, thanks to the added flexibility of learning the NBP.
@@ Acknowledgements
Mario Beraha gratefully acknowledges support from the Italian Ministry of Education, University and Research (MUR), “Dipartimenti di Eccellenza” grant 2023-2027. Jesper Møller is upported by The Danish Council for Independent Research — Natural Sciences, grant DFF – 10.46540/2032-00005B.

### v3

@@ Abstract
This paper proposes a novel approach for statistical modelling of a continuous random variable $X$ on $[0,1)$, based on its digit representation $X=.X_{1}X_{2}\ldots$. In general, $X$ can be coupled with a latent random variable $N$ so that $(X_{1},\ldots,X_{N})$ becomes a sufficient statistics and $X_{N+1}X_{N+2}\ldots$ is uniformly distributed. In line with this fact, and focusing on binary digits for simplicity, we propose a family of generalized finite Pólya trees that induces a random density for a sample, which becomes a flexible tool for density estimation. Here, the digit system may be random and learned from the data. We provide a detailed Bayesian analysis, including closed form expression for the posterior distribution. We analyse the frequentist properties as the sample size increases, and provide sufficient conditions for consistency of the posterior distributions of the random density and $N$. We consider an extension to data spanning multiple orders of magnitude, and propose a prior distribution that encodes the so-called extended Newcomb-Benford law. Such a model shows promising results for density estimation of human-activity data. Our methodology is illustrated on several synthetic and real datasets.
@@ 1.3 Outline
The Supplemental material collects the missing proofs in the present paper and, in continuation of our remarks in the last paragraph of Section 1.1, discusses the case of data that lies outside the unit interval, see also the summary at the beginning of Section 6.
@@ 6 Numerical illustrations
This section investigates several aspects of our models via numerical simulations as follows. Section 6.1 assesses the impact of assuming $N$ random on density estimation by comparing the GFPT1 and standard (finite) Pólya tree priors, focusing on accuracy of the density estimates as well as prior and posterior variability. Section 6.2 compares the GFPT1 and GFPT2 priors in terms of density estimation accuracy and the posterior distribution of $N$. Moreover, we compare our models with the Optional Pólya tree (OPT) model of Wong and Ma (2010) and the APT model of Ma (2017). Section 6.3 illustrates the advantages of the MBPT model on real datasets spanning multiple orders of magnitude. Finally, Appendix F in Supplemental material discusses the case of data that lies outside the unit interval, and how the use of a bijection to transform data living in $\mathbb{R}$ or $\mathbb{R}_{+}$ to $[0,1)$ affects density estimation, as well as the role of $N$ as a smoothing parameter. In particular, we observe that care must be taken when applying such a transformation to avoid that most of the datapoints get mapped near the boundaries of the interval $[0,1)$: A simple scaling of the data by their empirical standard deviation before applying the transformation yields accurate posterior density estimates. Moreover, the importance of $N$ is unchanged, as it allows to adapt to the smoothness of the data.
Prior elicitation follows the discussion in Section 4.3. Unless otherwise specified, a priori $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,20\}$ where the Poisson distribution has mean 5.
@@ 6.1 Posterior results when using GFPT1 and Pólya tree priors
Figure 4 show the mean across the 100 independent datasets of
@@
respectively. As expected Figure 4 show that larger values of $\alpha_{0}$ correspond to smaller values of $N$ (a posteriori) and vice versa. This is particularly evident for datasets DG1 and DG3. In general, we see that GFPT1 priors outperform PT priors in terms of the quality of the density estimate. Moreover, larger $\alpha_{0}$ clearly yield less wiggly functions when using PT priors, while this dependence is not so clear when using GFPT1 priors. In conclusion, we find that $\alpha_{0}=0.1$ provides the best trade off between wigglyness and accuracy of the density estimate, and we suggest using it as default value for GFPT1 priors.
Figure 4: Posterior results for the second simulation study in Section 6.1. From top to bottom: Total variation (TV) distance between true and estimated densities, $W$-function for the difference between true and estimated densities, and posterior expectation of $N$ as a function of the sample size $m$. The three columns correspond to the data examples DG1–DG3, respectively. Figures are averaged over 100 independent replicates. Solid and dashed lines refer to the use of GFPT1 and PT priors, respectively. Different colors refer to different values of $\alpha_{0}$.
@@ 6.3 Fitted models for real datasets spanning multiple orders of magnitude
Data for the first four dataset can be found at https://github.com/jasonlong/benfords-law, while data for the income dataset is publicly available from https://www.census.gov/data.html. Table 1 reports, for each dataset, the empirical range (maximum and minimum value of the data) and two diagnostics computed from the log-mantissa $U=\log_{10}(Z)-\lfloor\log_{10}(Z)\rfloor$: The Wasserstein-1 distance $W_{1}$ between the empirical distribution of $U$ and the $\mathrm{Unif}[0,1)$ law, and the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ to $\mathrm{Unif}[0,1)$. Under the Newcomb-Benford law, the distribution is scale invariant, i.e., the log-mantissa $U$ is $\mathrm{Unif}[0,1)$-distributed, hence, small values of $W_{1}$ and $d_{\mathrm{KS}}$ indicate that our MBPT model is well ssupported by the data. The datasets in Table 1 span four, three, eight, six, and seven orders of magnitude, respectively, making them particularly well suited to test the adequacy of the base-2 and base-10 multiscale Benford Pólya tree priors. In addition, the $W_{1}$ and $d_{\mathrm{KS}}$ statistics quantify how close each dataset is to the Benford benchmark: eurodist has the largest deviations (0.18 and 0.07), signalling weak scale-invariance; by contrast, census is extremely close to uniform (0.001 and 0.003), with income, GDP, and Twitter also showing relatively small deviations. This pattern anticipates our empirical findings, whereby the MBPT model delivers the best fit where the data spans a large number of orders of magnitude and $W_{1}$ and $d_{\mathrm{KS}}$ are close to zero.
@@
The posterior inferences using the different priors are compared in term of the widely applicable information criterion (WAIC), see Watanabe (2013), which is a consistent estimator of the out-of-sample error (Vehtari et al., 2017). Since the absolute value of the WAIC is irrelevant, we fix the DPM as a baseline prior and report in Table 2 the relative improvement of the WAIC for any model over the DPM (larger values of the relative improvement correspond to better model performance, while negative values indicate that the DPM achieves a better fit to the data). For datasets eurodist, GDP, and census, all priors seem to perform similarly. For the datasets Twitter and income datasets, the multiscale Benford Pólya tree priors clearly outperforms all other priors, where in particular, the $\mathrm{MBPT}_{10}$ prior shows a remarkable improvement of roughly 19% and 25% over the DPM prior, respectively. These results confirm the intuition obtained by looking at the diagnostics based on the log-mantissa and the span of the datasets discussed above. Moreover, in all scenarios, the WAIC when using the GFPT2 prior slightly improves as compared to the GFPT1 prior. Although not reported here, we have also tried fitting a variant of the MBPT models where the parameters $\alpha$ are not specified in accordance to the Newcomb-Benford law but according to the standard choice of (16). We noticed substantial deterioration in the WAICs across all datasets, and in particular for Twitter, for which this variant performs worse than the DPM baseline, and income, for which this variant performs on par to the baseline.
@@
Finally, we comment on the number of sufficient digits needed in the different datasets according to the GFPT and MBPT models. Table 3 reports the (estimated) a posteriori expectation of $N$ in the different models. Note that $N$ is typically small a posteriori, despite $n_{\max}$ being equal to six, four, seventeen, eight, and six in the different datasets respectively. Recall that the prior for $N$ is truncated at $n_{\max}$ for the $\mathrm{MBPT}_{10}$ prior and to $\lceil\log_{2}(10^{n_{\max}})\rceil$ for the other priors. Hence, the sufficient number of digits are effectively always less than $n_{\max}$. Except for the Twitter and income datasets, we observe that $N$ tends to be smaller a posteriori under the $\mathrm{MBPT}_{2}$ prior than under the GFPT1 and GFPT2 priors. This entails that the max-scaling of the data disaligns the data from the nested binary partitions. Indeed, using the GFPT2 prior, the posterior adaptively learns the NBP, leading to a smaller $N$ than under the GFPT1 prior that works with the fixed NBP. For all the different datasets, a posteriori $N$ tends to be smaller under the $\mathrm{MBPT}_{10}$ prior, however, the possible values taken by $X_{1:N}$ is $10^{N}$ under the $\mathrm{MBPT}_{10}$ prior and (approximately) $2^{N}$ under the other models. We report the corresponding (estimated) posterior expectations in Table 3. When using the $\mathrm{MBPT}_{10}$ prior a larger number of possible digits is required a posteriori than if using another of the priors, but compared with the $\mathrm{MBPT}_{2}$ prior, the difference may not be extreme, since the scale is coarser for $10^{N}$ than for $2^{N}$. On the other hand, there are several situations in which the $\mathrm{GPFT2}$ prior causes a posteriori a much lower number of possible digits than if using another of the priors, thanks to the added flexibility of learning the NBP.
@@ Acknowledgements
Mario Beraha gratefully acknowledges support from the Italian Ministry of Education, University and Research (MUR), “Dipartimenti di Eccellenza” grant 2023-2027. Jesper Møller is supported by The Danish Council for Independent Research — Natural Sciences, grant DFF – 10.46540/2032-00005B.
@@
Appendix E Simulations for data on $\mathbb{R}$ or $\mathbb{R}_{+}$
As discussed at the end of Section 1.1, our methodology can be applied to data supported on spaces larger than $[0,1)$ by means of suitable transformations. In this appendix, we investigate the impact of such transformations on posterior inference.
E.1 Data supported on $\mathbb{R}$
We generate $m=1000$ observations from a continuous distribution on $\mathbb{R}$, specifically
$Z_{1},\ldots,Z_{m}\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{3}t_{3}(-4,1)+\frac{2}{3}\mathcal{N}(3,1),$
where $t_{a}(\mu,s)$ denotes the Student $t$ distribution with $a$ degrees of freedom, location $\mu$, and scale $s$.
Data are mapped to $[0,1)$ via a bijective transformation $g$. In particular, we consider the logistic function $g(x)=1/(1+e^{-x})$ and the probit function $g(x)=\Phi(x)$, where $\Phi$ is the CDF of the standard normal distribution. We fit a standard Pólya tree and a GFPT1 model (with the default prior specification as in Sections 6.1 and 6.2, and $\alpha_{0}=0.1$) to the transformed data and obtain posterior samples of the random PDF on $[0,1)$ induced by the two models. These posterior samples are then mapped back to $\mathbb{R}$ via the usual change-of-variable formula, and the density estimate is obtained by taking the pointwise sample average of the transformed densities.
Figure E1: Posterior inference for the simulation in Appendix E.1. Top row: posterior density estimates under PT and GFPT1 models when data are transformed using the logistic function, without (left) and with (right) proper scaling. Bottom row: data transformed using the probit function.
The two leftmost panels of Figure E1 show the density estimates obtained using the logistic and probit transformations without scaling. The resulting fits are rather poor, especially under the probit transformation, in the left and right tails of the distribution. This behaviour is a consequence of the fact that the transformation shrinks tail observations towards 0 and $1$, making it difficult for both the PT and GFPT1 models, when combined with the standard NBP $I$, to capture the tail behaviour.
One possibility would be to adopt an alternative NBP that allocates smaller intervals near the boundaries 0 and $1$ and larger ones in the interior of the unit interval. A simpler solution is to account for the scale of the data by using a scaled version of the transformation, $g(x/s)$, where $s$ is the empirical standard deviation of the data; we refer to these as transformations with proper scaling. The two rightmost panels of Figure E1 display the density estimates obtained when the properly scaled transformations are used. Posterior density estimates improve substantially, especially under the GFPT1 model. As in Section 6.1, the PT-based estimates remain extremely wiggly, whereas the GFPT1 prior yields smoother density estimates that better adapt to the underlying data-generating distribution.
E.2 Data supported on $\mathbb{R}_{+}$.
Figure E2: Density estimates for the simulation in Appendix E.2.
We generate $m=1000$ observations from
$Z_{1},\ldots,Z_{m}\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{2}\mathrm{Ga}(2,2)+\frac{1}{2}\mathrm{Ga}(3,1;8),$
where $\mathrm{Ga}(a,b;\mu)$ denotes the gamma distribution with shape $a$, rate $b$, and shift $\mu$ (which defaults to zero for the standard gamma distribution), that is, $X\sim\mathrm{Ga}(a,b;\mu)$ iff $X-\mu\sim\mathrm{Ga}(a,b)$.
We considered several bijective transformations $g$ from $\mathbb{R}_{+}$ to $[0,1)$, but most of them led to unsatisfactory results, with density estimates failing to capture the right tail of the true data-generating distribution. As in Section E.1, this is due to the fact that any bijection from $\mathbb{R}_{+}$ to $[0,1)$ necessarily maps large values close to $1$, and the PT and GFPT1 models on the standard NBP $I$ are not well suited to modelling such compressed tails.
In this setting, we found a two-stage approach to work particularly well. First, data are mapped from $\mathbb{R}_{+}$ to $\mathbb{R}$ using the logarithmic transformation. The transformed data are then centred so that their sample mean is zero, and the logistic transformation with proper scaling is applied. As before, once the models are fitted on $[0,1)$, posterior samples of the random PDF are mapped back to $\mathbb{R}_{+}$ via the change-of-variable formula.
Figure E2 displays the density estimates obtained with the GFPT1 and standard PT models; the same qualitative comments as in Appendix E.1 apply.